The length of a rectangle is 4 inches more than its width. If the area of the rectangle is 285 square inches, what is its length?
19 in
41 in
8 in
15 in

Answer by lwsshak3(11628)

The length of a rectangle is 4 inches more than its width. If the area of the rectangle is 285 square inches, what is its length?
let x=width of rectangle
x+4=length of rectangle
width*length=area
x*(x+4)=285
x^2+4x=285
x^2+4x-285=0
(x+19)(x-15)=0
x=15
x+4=19
what is its length? 19 inches